construct cayley table (Z subscript (9) , circle times )and verify its an abelian group.
1
Expert's answer
2020-10-21T16:06:13-0400

"\begin{matrix}
 |& 0 & 1 & 2 &3 &4 &5 &6 & 7& 8\\
\hline
 0| & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 \\
1|& 1 & 2 & 3 & 4& 5 &6 & 7 & 8 & 0\\
2 | & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 0 & 1\\
3| &3 & 4 & 5 & 6 & 7 & 8 & 0 &1 & 2\\
4 |& 4 & 5 & 6 & 7 & 8 & 0 &1 & 2 & 3\\
5|&5 & 6 & 7 & 8 & 0 & 1 & 2 & 3 & 4\\
6 |& 6 & 7 & 8 & 0 & 1 & 2 & 3 & 4 & 5\\
7|&7 &8 & 0 & 1 & 2 & 3 & 4 & 5 & 6\\
8 |& 8 &0 &1 & 2 & 3 & 4 & 5 & 6 & 7 \\

\end{matrix}"

Cayley table of (Z"_{9},+)" .

The table is symmetric. Hence one can check a+b=b+a for all a,b"\in Z_{9}."

All the entries in the cayley table are from "0,1,\cdots,8". Hence closure follows.

The column and row of 0 shows that 0 is the identity.

Every row and column contains 0. Hence inverse exist.

Associativity hold true follows from the fact that addition modulo 9 (any n>0) is an associative operation. Associativity also follows by a laborious check from the cayley table.


Therefore (Z"_{9},+)" is an abelian group.